To extend a spring by 7 cm a job of 14 J is performed. Find the coefficient of elasticity of the system.


1
Expert's answer
2021-04-12T07:01:42-0400

"PE=\frac{kx^2}{2}\to k=\frac{2PE}{x^2}=\frac{2\cdot 14}{0.07^2}=5714\ (N/m)" . Answer